OpenStudy (anonymous):

A 4.0×104kg locomotive, with steel wheels, is traveling at 9.0m/s on steel rails when its engine and brakes both fail. The coefficient of friction is 1.7×10−3.

OpenStudy (anonymous):

how far will it go before it stops?

OpenStudy (anonymous):

any help would greatly be appreciated...

OpenStudy (anonymous):

UncleRhakus can you help

OpenStudy (unklerhaukus):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[ m=4.0\times10^4[\text{kg}]\\v=9.0 [\text{m/s}]\\\mu=1.7\times10^{−3}\]

OpenStudy (anonymous):

ok, as for the formula i think i'll be using two different ones. Can you confirm that?

OpenStudy (unklerhaukus):

Yeah what are the relevant forces?

OpenStudy (anonymous):

Can you elaborate on your question? Are you referring to the Newtons or m coefficient of friciton of [1.7 x 10^{3}\]

OpenStudy (unklerhaukus):

isn't it -3 ?

OpenStudy (anonymous):

yes sorry. i'm new to the equation thing

OpenStudy (anonymous):

forces will be drag, friction

OpenStudy (unklerhaukus):

type like this ``` \[1.7 \times 10^{-3}\] ```

OpenStudy (unklerhaukus):

what about gravity?

OpenStudy (anonymous):

ok yea that also. -g

OpenStudy (unklerhaukus):

and the normal force

OpenStudy (anonymous):

perpendicular plane, so weight also of \[\4.0[ x \10^{4}]\]

OpenStudy (anonymous):

so force would be weight x -g

OpenStudy (anonymous):

Force weight=mass*gravity ?

OpenStudy (unklerhaukus):

The weight force \[F_w=-mg\]if we take up to be positive

OpenStudy (unklerhaukus):

yeah and the normal force will be \[F_N=+mg\]

OpenStudy (unklerhaukus):

(assuming the train track is not on a hill)

OpenStudy (anonymous):

no its flat

OpenStudy (anonymous):

I don't think I'm putting together the concepts very well.

OpenStudy (unklerhaukus):

the friction force will be \[F_f=\mu N\] and will point in the opposite direction to the motion of the train

OpenStudy (anonymous):

ok, i'll try to crunch it out. thanks

OpenStudy (unklerhaukus):

What is the question asking for/?

OpenStudy (unklerhaukus):

ops \(F_N=N\)

OpenStudy (unklerhaukus):

maybe the the friction force makes more sense as \[F_f=-μN\]